# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 9.2. Ornstein–Uhlenbeck limit on the normal space
$|\mathcal{J}_{1}^{(2)}|\lesssim\frac{\big|\Xi_{\varepsilon}^{c}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}{\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (10.9)
@@
$\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}:\min_{i\neq j}|\xi_{i}-\xi_{j}|<(1+\eta)d_{\varepsilon}\}\big|\sim O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon}).$ (10.10)
Hence, the separation $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$ removes only a thin tubular neighbourhood of the diagonal band, of thickness $d_{\varepsilon}$, from the large cube of side length $\overline{L}_{\varepsilon}$, whose volume is of order $O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon})$. Therefore, we have
$\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|$
$=\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\}\big|-O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon})$
$\sim(2\overline{L}_{\varepsilon})^{Q}\Big(1-O\Big(\frac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\Big)\Big)\sim(2\overline{L}_{\varepsilon})^{Q},$ (10.11)
where in the last line we used $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$ and $\overline{L}_{\varepsilon}\sim\varepsilon^{-\frac{1}{2}+\eta}$ from (4.2). So the outside band region has almost full area, up to a relative error of order $\tfrac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\sim\varepsilon^{\frac{1}{2}-}$. For the complement of the set $\Xi_{\varepsilon}$ defined in (10.5), we have
@@
$\lesssim Q(\overline{L}_{\varepsilon})^{Q-1}(|\operatorname*{supp}g|+2\varepsilon^{-\frac{1}{2}+2\eta}).$ (10.12)
@@
$\frac{\big|\Xi_{\varepsilon}^{c}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}{\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}\lesssim\frac{|\operatorname*{supp}g|+2\varepsilon^{-\frac{1}{2}+2\eta}}{\overline{L}_{\varepsilon}}\lesssim\varepsilon^{\eta},$
@@
$\int\mathbf{1}_{\{\pi_{\varepsilon}^{T}(\phi)\in A\}}\rho_{\varepsilon}^{Q}(d\phi)=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}),$ (11.1)
where
$\mathcal{J}_{1}$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}\mathbf{1}_{\{\pi_{\varepsilon}^{T}(\phi)\in A\}}\rho_{\varepsilon}^{Q}(d\phi).$
Here $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as defined in (6.28) and (6.29). By using the coordinate expression $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ and Proposition 7.4, we write
$\mathcal{J}_{1}$ $=\overline{Z}_{\varepsilon}^{-1}\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)\mathbf{1}_{\{(\xi_{1},\dots,\xi_{Q})\in A\}}\cdot\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}\,d\xi_{1}\dots d\xi_{Q},$ (11.2)
where
@@
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)$ $=\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\mathbf{1}_{A_{\varepsilon}})+\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\mathbf{1}_{A_{\varepsilon}^{c}})$
$=1+O(\varepsilon^{0+}),$ (11.3)
uniformly in $\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}$. By following the same procedure, we can also write the partition function $\overline{Z}_{\varepsilon}$ in (11.2) as follows
$\overline{Z}_{\varepsilon}=\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots d\xi_{Q}\cdot(1+O(\varepsilon^{0+})).$ (11.4)
@@
$\mathcal{J}_{1}=\frac{\big|A\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.5)
@@
$\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|$
$\sim(2\overline{L}_{\varepsilon})^{Q}\Big(1-O\Big(\frac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\Big)\Big)\sim(2\overline{L}_{\varepsilon})^{Q},$ (11.6)
Therefore, the effect of imposing $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$ is negligible compared to the total volume of the simplex $\frac{(2\overline{L}_{\varepsilon})^{Q}}{Q!}$. Using (11.5) and (11.6), we obtain
$\mathcal{J}_{1}=\frac{\big|A\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.7)
@@
$\rho_{\varepsilon}^{Q}\{\pi_{\varepsilon}^{T}(\phi)\in A\}=\frac{\big|A\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.8)
@@
$\rho^{Q}_{\varepsilon}\big\{\pi^{T}_{j}(\phi)\in B\big\}=\frac{\big|\{\xi_{j}\in B\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}<\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.9)
@@
$S_{j}(x)=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}:\xi_{j}=x\}.$ (11.10)
@@
$\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}.$ (11.11)
@@
$\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!}.$ (11.12)
@@
$\big|\{\xi_{j}\in B\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|=\int_{B}|S_{j}(x)|dx=\int_{B}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!}dx.$
@@
$f_{j}(x)=\frac{Q!}{(2L_{\varepsilon})^{Q}}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!},\;\;-\overline{L}_{\varepsilon}<x<\overline{L}_{\varepsilon}.$
@@ Acknowledgements.
H.S. gratefully acknowledges financial support from an NSF grant (CAREER DMS-2044415). The work of P.S. is partially supported by NSF grants DMS-1811093, DMS-2154090, and a Simons Fellowship.

### v2

@@ 9.2. Ornstein–Uhlenbeck limit on the normal space
(10.9) $|\mathcal{J}_{1}^{(2)}|\lesssim\frac{\big|\Xi_{\varepsilon}^{c}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|}{\big|\Delta_{Q}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$
@@
(10.10) $\big|\{(\xi_{1},\dots,\xi_{Q})\in\Delta_{Q}:\min_{i\neq j}|\xi_{i}-\xi_{j}|<d_{\varepsilon}\}\big|\sim O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon}).$
Hence, the separation $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$ removes only a thin tubular neighbourhood of the diagonal band, of thickness $d_{\varepsilon}$, from the large cube of side length $\overline{L}_{\varepsilon}$, whose volume is of order $O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon})$. Therefore, we have
$\big|\Delta_{Q}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|$ $=\big|\Delta_{Q}\big|-O\big(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon}\big)$
(10.11) $\sim(2\overline{L}_{\varepsilon})^{Q}\Big(1-O\Big(\frac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\Big)\Big)\sim(2\overline{L}_{\varepsilon})^{Q},$
where in the last line we used $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$ and $\overline{L}_{\varepsilon}\sim\varepsilon^{-\frac{1}{2}+\eta}$ from (4.2). So the outside band region has almost full area, up to a relative error of order $\tfrac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\sim\varepsilon^{\frac{1}{2}-}$. For the complement of the set $\Xi_{\varepsilon}$ defined in (10.5), we have
@@
(10.12) $\lesssim Q(\overline{L}_{\varepsilon})^{Q-1}(|\operatorname*{supp}g|+2\varepsilon^{-\frac{1}{2}+2\eta}).$
@@
$\frac{\big|\Xi_{\varepsilon}^{c}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|}{\big|\Delta_{Q}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|}\lesssim\frac{|\operatorname*{supp}g|+2\varepsilon^{-\frac{1}{2}+2\eta}}{\overline{L}_{\varepsilon}}\lesssim\varepsilon^{\eta},$
@@
(11.1) $\int\mathbf{1}_{\{\pi_{\varepsilon}^{T}(\phi)\in A\}}\rho_{\varepsilon}^{Q}(d\phi)=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}),$
where $\mathcal{J}_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}\mathbf{1}_{\{\pi_{\varepsilon}^{T}(\phi)\in A\}}\rho_{\varepsilon}^{Q}(d\phi)$. Here $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$, as defined in (6.29) and (6.30). By using the coordinate expression $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ and Proposition 7.2, we write
(11.2) $\mathcal{J}_{1}$ $=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Delta_{Q}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)\mathbf{1}_{\{(\xi_{1},\dots,\xi_{Q})\in A\}}\cdot\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}}\,d\xi_{1}\dots d\xi_{Q},$
where $\Delta_{Q}=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}$ is defined as in (4.17) and
@@
(11.3) $\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)$ $=\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\mathbf{1}_{A_{\varepsilon}})+\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\mathbf{1}_{A_{\varepsilon}^{c}})=1+O(\varepsilon^{0+}),$
uniformly in $(\xi_{1},\dots,\xi_{Q})\in\Delta_{Q}$. By following the same procedure, we can also write the partition function $\overline{Z}_{\varepsilon}$ in (11.2) as follows
(11.4) $\overline{Z}_{\varepsilon}=\int\cdots\int_{\Delta_{Q}}\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}}d\xi_{1}\dots d\xi_{Q}\cdot(1+O(\varepsilon^{0+})).$
@@
(11.5) $\mathcal{J}_{1}=\frac{\big|A\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\cap\Delta_{Q}\big|}{\big|\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\cap\Delta_{Q}\big|}(1+O(\varepsilon^{0+})).$
@@
(11.6) $\big|\Delta_{Q}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|$ $\sim(2\overline{L}_{\varepsilon})^{Q}\Big(1-O\Big(\frac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\Big)\Big)\sim(2\overline{L}_{\varepsilon})^{Q}.$
Therefore, the effect of imposing $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$ is negligible compared to the total volume of the simplex $\frac{(2\overline{L}_{\varepsilon})^{Q}}{Q!}$. Using (11.5) and (11.6), we obtain
(11.7) $\mathcal{J}_{1}=\frac{|A\cap\Delta_{Q}|}{|\Delta_{Q}|}(1+O(\varepsilon^{0+})).$
@@
(11.8) $\rho_{\varepsilon}^{Q}\{\pi_{\varepsilon}^{T}(\phi)\in A\}=\frac{|A\cap\Delta_{Q}|}{|\Delta_{Q}|}(1+O(\varepsilon^{0+})).$
@@
(11.9) $\rho^{Q}_{\varepsilon}\big\{\pi^{T}_{j}(\phi)\in B\big\}=\frac{|\{\xi_{j}\in B\}\cap\Delta_{Q}|}{|\Delta_{Q}|}(1+O(\varepsilon^{0+})).$
@@
(11.10) $S_{j}(x)=\{(\xi_{1},\dots,\xi_{Q})\in\Delta_{Q}:\xi_{j}=x\}.$
@@
(11.11) $\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}.$
@@
(11.12) $\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!}.$
@@
$|\{\xi_{j}\in B\}\cap\Delta_{Q}|=\int_{B}|S_{j}(x)|dx=\int_{B}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!}dx.$
@@
$f_{j}(x)=\frac{Q!}{(2\overline{L}_{\varepsilon})^{Q}}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!},\;\;-\overline{L}_{\varepsilon}<x<\overline{L}_{\varepsilon}.$
@@ Acknowledgements.
K.S. would like to thank Hendrik Weber for pointing out helpful references. H.S. gratefully acknowledges financial support from an NSF grant (CAREER DMS-2044415). The work of P.S. is partially supported by NSF grants DMS-1811093, DMS-2154090, and a Simons Fellowship.